# SemantiCodec: An Ultra Low Bitrate Semantic Audio Codec for General Sound

### v1

@@ V-C2 Acoustic Representation Learning
The size of the acoustic codebook determines the granularity of the trainable vector quantization module. Analogous to the influence of k-means centroid numbers, larger codebook size leads to better reconstruction quality, as indicated by Table V.
@@
TABLE V: The impact of SemantiCodec encoder parameters on the reconstruction quality.
@@ V-C3 Learnable Semantic Codebook
TABLE VI: The impact of replacing the k-means centroid with a learnable vector quantization layer.
@@
Learnable Semantic VQ $3.57$ $5.1$ $28.5$
The semantic codebook in SemantiCodec is frozen during model training. To validate the importance of performing k-means clustering beforehand, we conduct another experiment by replacing the semantic codebook with a learnable VQ layer, similar to the acoustic VQ layer. As shown in Table VI, even though the reconstruction performance of SemantiCodec is not largely affected without the use of k-means centroids, the semantic evaluation accuracy shows a notable decrease from $53.1$% to $28.5$%. This indicates the validity of k-means centroids used in the semantic codebook for maintaining rich semantic information in audio tokens. Directly training a codec model with two VQ layers results in the codec neglecting semantic information and focusing solely on reconstruction capabilities, as the training loss function is limited to reconstruction loss.
@@ V-C4 DDIM Sampling Setups
The sampling step also plays a crucial role in the reconstruction quality of our latent diffusion decoder. The reconstruction quality of SemantiCodec is moderate, with a small number of sampling steps, e.g., ten steps. With a sampling step larger than $25$, the reconstruction quality significantly improves. This is promising for rapid sampling during audio reconstruction with tokens from SemantiCodec.
@@ VI Conclusion
In this study, we have proposed SemantiCodec, an audio codec which can be applied to diverse audio types with ultra-low bitrates and rich semantic information in the audio tokens. SemantiCodec supports several bitrates between $0.31$kbps and $1.43$kbps. With a semantic and acoustic decoupled architecture, SemantiCodec achieves effective compression without significantly sacrificing quality with a token rate of $50$ per second and $100$ per second. At an ultra-low rate of $25$ tokens per second, SemantiCodec still demonstrates significantly strong audio quality compared with state-of-the-art audio codecs. Our experiment result also shows that the semantic information within the SemantiCodec tokens is significantly richer than previous neural audio codecs.
@@ Acknowledgments
This research was partly supported by the British Broadcasting Corporation Research and Development (BBC R&D), Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound”, and a PhD scholarship from the Centre for Vision, Speech and Signal Processing (CVSSP), Faculty of Engineering and Physical Science (FEPS), University of Surrey. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.

### v2

@@ V-C2 Acoustic Representation Learning
The size of the acoustic codebook determines the granularity of the trainable vector quantization module. Analogous to the influence of k-means centroid numbers, larger codebook size leads to better reconstruction quality, as indicated by Table VI.
@@
TABLE VI: The impact of SemantiCodec encoder parameters on the reconstruction quality.
@@ V-C3 Learnable Semantic Codebook
TABLE VII: The impact of replacing the k-means centroid with a learnable vector quantization layer.
@@
Learnable Semantic VQ $3.57$ $5.10$ $28.5$
The semantic codebook in SemantiCodec is frozen during model training. To validate the importance of performing k-means clustering beforehand, we conduct another experiment by replacing the semantic codebook with a learnable VQ layer, similar to the acoustic VQ layer. As shown in Table VII, even though the reconstruction performance of SemantiCodec is not largely affected without the use of k-means centroids, the semantic evaluation accuracy shows a notable decrease from $53.1$% to $28.5$%. This indicates the validity of k-means centroids used in the semantic codebook for maintaining rich semantic information in audio tokens. Directly training a codec model with two VQ layers results in the codec neglecting semantic information and focusing solely on reconstruction capabilities, as the training loss function is limited to reconstruction loss.
@@ V-C4 DDIM Sampling Setups
The sampling step also plays a crucial role in the reconstruction quality of our latent diffusion decoder. The reconstruction quality of SemantiCodec is moderate, with a small number of sampling steps, e.g., ten steps. With a number of sampling steps larger than $25$, the reconstruction quality significantly improves. This is promising for rapid sampling during audio reconstruction with tokens from SemantiCodec.
@@ VI Conclusion
In this study, we have proposed SemantiCodec, an audio codec which can be applied to diverse audio types with ultra-low bitrates and rich semantic information in the audio tokens. SemantiCodec supports several bitrates between $0.31$ kbps and $1.40$ kbps. With a semantic and acoustic decoupled architecture, SemantiCodec achieves effective compression without significantly sacrificing quality with token rates of $50$ and $100$ per second. At an ultra-low rate of $25$ tokens per second, SemantiCodec still demonstrates significantly strong audio quality compared with state-of-the-art audio codecs. Our experiment result also shows that the semantic information within the SemantiCodec tokens is significantly richer than in previous neural audio codecs.
@@ Acknowledgments
This research was partly supported by the British Broadcasting Corporation Research and Development (BBC R&D), Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound”, and a PhD scholarship from the Centre for Vision, Speech and Signal Processing (CVSSP), Faculty of Engineering and Physical Science (FEPS), University of Surrey. This publication is supported by multiple datasets that are openly available at locations referenced in this paper. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising. We would also like to express our gratitude to the anonymous reviewers for their valuable comments.